# supplementary

### v1

@@ 3.3 Connection to direct regularization methods
Equi–sparsity can also be achieved by pairwise fusion, e.g., using the fused LASSO (Tibshirani et al., 2005), which minimizes a loss plus $\sum_{i<j}w_{ij}\lvert\beta_{i}-\beta_{j}\rvert$. Related types of pairwise fusion include the 1-D segmented version (Tibshirani et al., 2005), graph-guided fused LASSO (Chen et al., 2010), OSCAR’s $L_{\infty}$ variant (Bondell and Reich, 2008), and several adaptively weighted variants (Ke et al., 2015; Tang and Song, 2016; Chu et al., 2021). They’re further related to the topic of convex clustering(Chi et al., 2025). These methods, however, lack a built-in mechanism for enforcing tree-coherence in our motivating context. Moreover, their theoretical guarantees often require an accurate pilot estimator.
Our penalty generalizes fusion: for any pair $\{\beta_{1},\beta_{2}\}$, $\lvert\beta_{1}-\beta_{2}\rvert=\sqrt{2}\,\bigl{\|}D_{2}(\beta_{1},\beta_{2})^{\!\top}\bigr{\|}_{2},$ so the fused LASSO term is the $a_{\ell}=2$ instance of our group penalty. Allowing $a_{\ell}>2$ enforces group-wise fusion, naturally accommodates the tree structure and also makes the approach widely applicable to other aggregation tasks.
@@ 4.1 Overview
The optimization problem (3.1) combines a smooth loss $g$ and the nonsmooth but convex tree penalty $\lambda\Omega_{\mathcal{T},w}$. For linear and logistic models we use the usual quadratic and cross-entropy losses
@@
[…]
@@ 4.2 Proximal operator for $\Omega_{\mathcal{T},w}$
The proximal map
$\operatorname{prox}_{\lambda\Omega_{\mathcal{T},w}}(\eta)=\mathop{\rm argmin}\limits_{\beta}\Bigl{\{}\tfrac{1}{2}\|\beta-\eta\|_{2}^{2}+\lambda\Omega_{\mathcal{T},w}(\beta)\Bigr{\}}$ (4.2)
is uniquely defined because the objective is strongly convex. Although the tree‐structured penalty precludes a closed-form solution in general, the map can be evaluated exactly by a “bottom-up” sequence of shrinkage updates, each of which has a simple closed form. To illustrate this algorithm, define
$\tilde{\beta}_{0}=\mathop{\rm argmin}\limits_{\beta}\Bigl{\{}\tfrac{1}{2}\|\beta-\eta\|_{2}^{2}+\lambda\!\!\sum_{\ell\in\mathcal{I}^{\star}}\!\!w_{\ell}\|D_{a_{\ell}}\beta_{\mathcal{A}_{\ell}}\|_{2}\Bigr{\}},$ (4.3)
@@
where $\{\mathcal{A}_{\ell}\}_{\ell\in\mathcal{I}^{*}}$ and $\mathcal{A}_{\ell_{0}}$ correspond to internal nodes $\{b_{\ell}\}_{\ell\in\mathcal{I}^{*}}\subset\mathcal{T}$ and $b_{\ell_{0}}$. Going from (4.3) to (4.4), we add the penalty on group $\mathcal{A}_{\ell_{0}}$. In the following, we assume that corresponding internal node $b_{\ell_{0}}$ is not a descendant of any node in $\{b_{\ell}\}_{\ell\in\mathcal{I}^{\star}}$. In this case, the following theorem shows that $\hat{\beta}$ can be expressed in closed-form using $\hat{\beta}_{0}$.
@@
$[\tilde{\beta}]_{\mathcal{A}_{\ell_{0}}}$ =$=\begin{cases}\bar{\beta}_{\ell_{0}}\mathbf{1}_{a_{\ell_{0}}},&\rho_{\ell_{0}}\geq 1\\[4.0pt] \rho_{\ell_{0}}\bar{\beta}_{\ell_{0}}\mathbf{1}_{a_{\ell_{0}}}+(1-\rho_{\ell_{0}})[\tilde{\beta}_{0}]_{\mathcal{A}_{\ell_{0}}},&\rho_{\ell_{0}}<1\end{cases}~~~\text{and }~~[\tilde{\beta}]_{\mathcal{A}^{c}_{\ell_{0}}}=[\tilde{\beta}_{0}]_{\mathcal{A}^{c}_{\ell_{0}}},$
where $\mathcal{A}_{\ell_{0}}^{c}=[p]\setminus\mathcal{A}_{\ell_{0}}$ , and $[\tilde{\beta}]_{\mathcal{A}_{\ell_{0}}}$ is the subvector of $\tilde{\beta}$ indexed by $\mathcal{A}_{\ell_{0}}$.
@@
Applying the result of Theorem 4.1, we see that the proximal operator (4.2) can be computed using a bottom-up, one pass algorithm wherein we traverse the tree from leaves to root and sequentially add penalty terms. To start with, we group the latent nodes by depth: $(\mathcal{J}_{h},\ldots,\mathcal{J}_{0})$, where $\mathcal{J}_{0}$ contains only the roots. We describe the procedure in Algorithm 1, which is also visualized in Figure 2.
@@
5:   Solve it by applying Theorem  4.1
6:  end for
7:  return $\hat{\beta}^{(h)}$
8:end procedure
@@
$\mathbf{1}_{a_{\ell_{0}}}^{\top}[\tilde{\beta}]_{\mathcal{A}_{\ell_{0}}}=\mathbf{1}_{a_{\ell_{0}}}^{\top}[\tilde{\beta}_{0}]_{\mathcal{A}_{\ell_{0}}},\qquad\mathbf{1}_{p-a_{\ell_{0}}}^{\top}[\tilde{\beta}]_{\mathcal{A}^{c}_{\ell_{0}}}=\mathbf{1}_{p-a_{\ell_{0}}}^{\top}[\tilde{\beta}_{0}]_{\mathcal{A}^{c}_{\ell_{0}}}.$
@@ 5.1 Preliminaries
$\hat{\beta}\in\mathop{\rm argmin}_{\beta\in\mathbb{R}^{p}}\left\{g(\beta)+\lambda\Omega_{\mathcal{T},w}(\beta)\right\}$
@@
Throughout the section, we use default choice of weight $w_{\ell}=a_{\ell}^{-1/2}$, so that the penalty can be expressed $\Omega_{\mathcal{T},w}(\beta)=\sum_{\ell\in\mathcal{I}}\frac{1}{\sqrt{a_{\ell}}}\|D_{a_{\ell}}\beta_{\mathcal{A}_{\ell}}\|_{2}.$ To make the analysis easier, we define the matrix $M(\mathcal{A}_{\ell})\in\mathbb{R}^{p\times p}$ such that the submatrix with rows and columns indexed by $\mathcal{A}_{\ell}$ being equal to $D_{a_{\ell}}$ with the other entries being zero. As such, $[M(\mathcal{A}_{\ell})\beta]_{\mathcal{A}_{\ell}}=D_{a_{\ell}}\beta_{\mathcal{A}_{\ell}}$, and the penalty can be equivalently written as $\Omega_{\mathcal{T},w}(\beta)=\sum_{\ell\in[L]}\frac{1}{\sqrt{a_{\ell}}}\|M(\mathcal{A}_{\ell})\beta\|_{2}.$ Note that when $\mathcal{A}_{\ell}$ is a singleton representing a leaf node, $a_{\ell}=1$ means $D_{a_{\ell}}=0$ and $M(\mathcal{A}_{\ell})=(\mathbf{0})_{p\times p}$ is a zero matrix, then the corresponding penalty term is zero.
@@
The kernel consists of the directions along which $\Omega_{\mathcal{T},\omega}(\beta)$ stays unchanged when $\beta$ is perturbed. Intuitively, we can view the kernel node set as the roots of the tree (or the forest) structure. We take the example tree in Figure 1(a) for illustration. The penalty $\Omega_{\mathcal{T},w}(\beta)=7^{-1/2}\|M(\mathcal{A}_{{11}})\beta\|_{2}+3^{-1/2}\|M(\mathcal{A}_{9})\beta\|_{2}+2^{-1/2}\|M(\mathcal{A}_{8})\beta\|_{2}+2^{-1/2}\|M(\mathcal{A}_{{10}})\beta\|_{2}$ shrinks parameter $\beta$ towards the subspace $\operatorname{Ker}(\Omega_{\mathcal{T}},\omega)=\mathop{\rm span}\{\mathbf{1}_{7}\}$, which corresponds to the kernel node set $\mathcal{N}=\{11\}$, a singleton of the top layer node. We further study the property of the kernel in Appendix.
@@ 5.2 Condition-free prediction error bound for linear regression
We first consider the linear regression setting with data generating model $y=X\beta^{*}+\epsilon,$ and treat the design matrix as deterministic to study the predictive error $\|X\hat{\beta}-X\beta^{*}\|_{2}^{2}$. In this section, we impose only minimal regularity conditions on $X$. Thus, we call our error bound a “condition-free error bound”. We will make the following assumptions.
@@
[…]
@@
$\lambda$ $=\frac{4\sqrt{2}\sigma C}{\sqrt{n}}\Theta(\mathcal{T})\sqrt{\log(2p)+\log(|\mathcal{I}|)}~~\text{ where }~~~\Theta(\mathcal{T})=\sqrt{\max_{\ell\in\mathcal{I}}a_{\ell}}\left(1+\sqrt{\frac{\max_{\ell\in\mathcal{I}}{a_{\ell}}}{\log(|\mathcal{I}|)}}\right),$
then with probability at least $1-2/p$,
$\begin{split}\frac{1}{n}\|X(\hat{\beta}-\beta^{*})\|_{2}^{2}&\leq\frac{24\sqrt{2}\sigma C}{\sqrt{n}}\Theta(\mathcal{T})\sqrt{\log(2p)+\log(|\mathcal{I}|)}\left(\sum_{\ell_{1}\in\mathcal{I}_{1}}\frac{1}{\sqrt{a_{\ell_{1}}}}\|M(\mathcal{A}_{\ell_{1}})\beta^{*}\|_{2}\right)\\ &\quad\quad+\frac{32\sigma^{2}\{\log(2p)+\min(|\mathcal{N}|,n)\log(5)\}}{n},\end{split}$ (5.1)
@@
The quantity $\Theta(\mathcal{T})$ is defined regarding the given tree. Intuitively, $\Theta(\mathcal{T})$ grows with the size of the largest group and the number of nodes indicated by the tree. It quantifies the selection complexity of the group-wise penalty analogous to the analysis of group LASSO(Wainwright, 2019)[Chapter 9].
Unlike typical high-dimensional analysis, the result in Theorem 5.1 features two additive terms. The second term arises from the penalty’s null space whose cardinality $|\mathcal{N}|$, understood as the number of roots in the tree (or the forest) structure, is typically much smaller than $n$ and $p$. Hence, the second term is square order for $\sqrt{\log(p)/n}$ and usually becomes negligible with large $n$. Consequently, the first term is more influential in determining the primary error rate. Its main driver is
@@ 5.2.1 Direct Comparison with RARE
Consider a setting where the number of true groups in $\beta^{*}$ grows linearly with dimension $p$, so neither $K$ nor $p$ is considered constant in the analysis. In this situation, the tree is made up of $t$ copies of the same subtree, each containing $p_{s}$ parameters within the $K_{s}$ groups. Taking $p_{s}$ and $K_{s}$ as constants, the total dimension $p=t\,p_{s}$ and the total number of groups $K=tK_{s}$ increase with the number of subtree copies $t$. The following corollary shows that, under this scenario, the resulting predictive error rate is at least as good as LASSO’s “slow rate,” matching the rate of Yan and Bien (2021).
@@
[…]
@@
[…]
The result of the preceding corollary suggests an improvement by a factor of $1/\sqrt{\log(p)}$ over the method of Yan and Bien (2021), which is further verified by simulation results.

### v2

@@ 3.3 Connection to direct regularization methods
Equi–sparsity can also be achieved by pairwise fusion, e.g., using the fused LASSO (Tibshirani et al., 2005), which minimizes a loss plus $\sum_{i<j}w_{ij}\lvert\beta_{i}-\beta_{j}\rvert$. Related types of pairwise fusion include the 1-D segmented version (Tibshirani et al., 2005), graph-guided fused LASSO (Chen et al., 2010), OSCAR’s $L_{\infty}$ variant (Bondell and Reich, 2008), and several adaptively weighted variants (Ke et al., 2015; Tang and Song, 2016; Chu et al., 2021). These fusion methods are further related to the topic of convex clustering (Chi et al., 2025). These methods, however, lack a built-in mechanism for enforcing tree-coherence in our motivating context. Moreover, their theoretical guarantees often require an accurate pilot estimator.
Also worth noting is that our penalty generalizes fusion: for any pair $\{\beta_{1},\beta_{2}\}$, $\lvert\beta_{1}-\beta_{2}\rvert=\sqrt{2}\,\bigl\|D_{2}(\beta_{1},\beta_{2})^{\!\top}\bigr\|_{2},$ so the fused LASSO term is the $a_{\ell}=2$ instance of our group penalty. Allowing $a_{\ell}>2$ enforces group-wise fusion, naturally accommodates the tree structure and also makes the approach widely applicable to other aggregation tasks.
@@ 4.1 Overview
The optimization problem (3.1) combines a smooth loss $g$ and the nonsmooth but convex tree penalty $\lambda\Omega_{\mathcal{T},w}$. For linear and logistic regression, we use the usual quadratic and cross-entropy losses
@@
In both cases, $g$ is convex and differentiable with Lipschitz-continuous gradient. Thus, we can apply the accelerated proximal gradient descent algorithm (Nesterov, 1983; Parikh and Boyd, 2014), with details provided in the Supplementary Materials LABEL:appendix.compute. The algorithm iteratively updates the optimization variable $\beta$ through the proximal operator of $\lambda\Omega_{\mathcal{T},w}$. Specifically, for a step size $\tau>0$, the $(t+1)$-th iterate is given by
@@
Code implementing our method, along with code to reproduce all the results in this article, is available at https://github.com/JinwenFu001/direct_feature_aggregation.
@@ 4.2 Proximal operator for $\Omega_{\mathcal{T},w}$
The proximal mapping
$\vskip-15.0pt\operatorname{prox}_{\lambda\Omega_{\mathcal{T},w}}(\eta)=\mathop{\rm argmin}\limits_{\beta}\Bigl\{\tfrac{1}{2}\|\beta-\eta\|_{2}^{2}+\lambda\Omega_{\mathcal{T},w}(\beta)\Bigr\}$ (4.2)
is uniquely defined because the objective in (4.2) is strongly convex. Although the tree‐structured penalty precludes a closed-form solution in general, the map can be evaluated exactly by a “bottom-up” sequence of shrinkage updates, each of which has a simple closed form. To illustrate this algorithm, define
$\vskip-15.0pt\tilde{\beta}_{0}=\mathop{\rm argmin}\limits_{\beta}\Bigl\{\tfrac{1}{2}\|\beta-\eta\|_{2}^{2}+\lambda\!\!\sum_{\ell\in\mathcal{I}^{\star}}\!\!w_{\ell}\|D_{a_{\ell}}\beta_{\mathcal{A}_{\ell}}\|_{2}\Bigr\},$ (4.3)
@@
where $\{\mathcal{A}_{\ell}\}_{\ell\in\mathcal{I}^{*}}$ and $\mathcal{A}_{\ell_{0}}$ correspond to internal nodes $\{b_{\ell}\}_{\ell\in\mathcal{I}^{*}}\subset\mathcal{T}$ and $b_{\ell_{0}}$. Going from (4.3) to (4.4), we add the penalty on group $\mathcal{A}_{\ell_{0}}$. In the following, we assume that corresponding internal node $b_{\ell_{0}}$ is not a descendant of any node in $\{b_{\ell}\}_{\ell\in\mathcal{I}^{\star}}$. In this case, the following theorem shows that $\tilde{\beta}$ can be expressed in closed form using $\tilde{\beta}_{0}$.
@@
$[\tilde{\beta}]_{\mathcal{A}_{\ell_{0}}}$ $=\begin{cases}\bar{\beta}_{\ell_{0}}\mathbf{1}_{a_{\ell_{0}}},&\rho_{\ell_{0}}\geq 1\\[4.0pt] \rho_{\ell_{0}}\bar{\beta}_{\ell_{0}}\mathbf{1}_{a_{\ell_{0}}}+(1-\rho_{\ell_{0}})[\tilde{\beta}_{0}]_{\mathcal{A}_{\ell_{0}}},&\rho_{\ell_{0}}<1\end{cases}~~~\text{and }~~[\tilde{\beta}]_{\mathcal{A}^{c}_{\ell_{0}}}=[\tilde{\beta}_{0}]_{\mathcal{A}^{c}_{\ell_{0}}},$
where $\mathcal{A}_{\ell_{0}}^{c}=[p]\setminus\mathcal{A}_{\ell_{0}}$ , and $\tilde{\beta}]_{\mathcal{A}_{\ell_{0}}}$ is the subvector of $\tilde{\beta}$ indexed by $\mathcal{A}_{\ell_{0}}$.
@@
Applying the result of Theorem 4.1, we see that the proximal operator (4.2) can be computed using a bottom-up, one-pass algorithm wherein we traverse the tree from leaves to root and sequentially add penalty terms. To start with, we group the latent nodes by depth: $(\mathcal{J}_{h},\ldots,\mathcal{J}_{0})$, where $\mathcal{J}_{0}$ contains only the roots. We describe the procedure in Algorithm 1, which is also visualized in Figure 2.
@@
5:  end for
6:  return $\hat{\beta}^{(h)}$
7:end procedure
@@
$\vskip-15.0pt\mathbf{1}_{a_{\ell_{0}}}^{\top}[\tilde{\beta}]_{\mathcal{A}_{\ell_{0}}}=\mathbf{1}_{a_{\ell_{0}}}^{\top}[\tilde{\beta}_{0}]_{\mathcal{A}_{\ell_{0}}},\qquad\mathbf{1}_{p-a_{\ell_{0}}}^{\top}[\tilde{\beta}]_{\mathcal{A}^{c}_{\ell_{0}}}=\mathbf{1}_{p-a_{\ell_{0}}}^{\top}[\tilde{\beta}_{0}]_{\mathcal{A}^{c}_{\ell_{0}}}.$
@@ 5.1 Preliminaries
$\vskip-10.0pt\hat{\beta}\in\mathop{\rm argmin}_{\beta\in\mathbb{R}^{p}}\left\{g(\beta)+\lambda\Omega_{\mathcal{T},w}(\beta)\right\}$
@@
Throughout the section, we use default choice of weight $w_{\ell}=a_{\ell}^{-1/2}$, so that the penalty can be expressed $\Omega_{\mathcal{T},w}(\beta)=\sum_{\ell\in\mathcal{I}}\frac{1}{\sqrt{a_{\ell}}}\|D_{a_{\ell}}\beta_{\mathcal{A}_{\ell}}\|_{2}.$ To make the analysis easier, we define the matrix $M(\mathcal{A}_{\ell})\in\mathbb{R}^{p\times p}$ such that its submatrix with rows and columns indexed by $\mathcal{A}_{\ell}$ being equal to $D_{a_{\ell}}$ with the other entries being zero. As such, $M(\mathcal{A}_{\ell})\beta]_{\mathcal{A}_{\ell}}=D_{a_{\ell}}\beta_{\mathcal{A}_{\ell}}$, and the penalty can be equivalently written as $\Omega_{\mathcal{T},w}(\beta)=\sum_{\ell\in[L]}\frac{1}{\sqrt{a_{\ell}}}\|M(\mathcal{A}_{\ell})\beta\|_{2}.$ Note that when $\mathcal{A}_{\ell}$ is a singleton representing a leaf node, $a_{\ell}=1$ means $D_{a_{\ell}}=0$ and $M(\mathcal{A}_{\ell})=(\mathbf{0})_{p\times p}$ is a zero matrix, then the corresponding penalty term is zero.
@@
The kernel consists of the directions along which $\Omega_{\mathcal{T},\omega}(\beta)$ stays unchanged when $\beta$ is perturbed. Intuitively, we can view the kernel node set as the roots of the tree (or the forest) structure. We take the example tree in Figure 1(b) for illustration. The penalty $\Omega_{\mathcal{T},w}(\beta)=7^{-1/2}\|M(\mathcal{A}_{{11}})\beta\|_{2}+3^{-1/2}\|M(\mathcal{A}_{9})\beta\|_{2}+2^{-1/2}\|M(\mathcal{A}_{8})\beta\|_{2}+2^{-1/2}\|M(\mathcal{A}_{{10}})\beta\|_{2}$ shrinks parameter $\beta$ towards the subspace $\operatorname{Ker}(\Omega_{\mathcal{T}},\omega)=\mathop{\rm span}\{\mathbf{1}_{7}\}$, which corresponds to the kernel node set $\mathcal{N}=\{11\}$, a singleton of the top layer node. We further study the property of the kernel in the Supplementary Materials LABEL:appendix.dimension.
@@ 5.2 Condition-free prediction error bound for linear regression
We first consider the linear regression setting with data generating model $y=X\beta^{*}+\epsilon,$ and treat the design matrix as deterministic to study the predictive error $\|X\hat{\beta}-X\beta^{*}\|_{2}^{2}$. In this section, we impose only minimal regularity conditions on $X$. Thus, we call our error bound a “condition-free” error bound. We will make the following assumptions.
@@
There exists a constant $C$, such that $\max_{\ell\in\mathcal{I}}\lvert\!\lvert\!\lvert\frac{X_{\cdot\mathcal{A}_{\ell}}}{\sqrt{n}}\rvert\!\rvert\!\rvert\leq C<\infty$, where $\lvert\!\lvert\!\lvert\cdot\rvert\!\rvert\!\rvert$ is the spectral norm of its matrix-valued input.
@@
$\vskip-10.0pt\lambda=\frac{4\sqrt{2}\sigma C}{\sqrt{n}}\Theta(\mathcal{T})\sqrt{\log(2p)+\log(|\mathcal{I}|)}~~\text{ where }~~~\Theta(\mathcal{T})=\sqrt{\max_{\ell\in\mathcal{I}}a_{\ell}}\left(1+\sqrt{\frac{\max_{\ell\in\mathcal{I}}{a_{\ell}}}{\log(|\mathcal{I}|)}}\right),$
then with probability at least 12/p1-2/p,
$\begin{split}\frac{1}{n}\|X(\hat{\beta}-\beta^{*})\|_{2}^{2}&\leq\frac{24\sqrt{2}\sigma C}{\sqrt{n}}\Theta(\mathcal{T})\sqrt{\log(2p)+\log(|\mathcal{I}|)}\left(\sum_{\ell_{1}\in\mathcal{I}_{1}}\frac{1}{\sqrt{a_{\ell_{1}}}}\|M(\mathcal{A}_{\ell_{1}})\beta^{*}\|_{2}\right)\\ &\quad\quad+\frac{32\sigma^{2}\{\log(2p)+\min(|\mathcal{N}|,n)\log(5)\}}{n},\vskip-10.0pt\end{split}$ (5.1)
@@
The quantity $\Theta(\mathcal{T})$ is defined regarding the given tree. Intuitively, $\Theta(\mathcal{T})$ grows with the size of the largest group and the number of nodes indicated by the tree. It quantifies the selection complexity of the group-wise penalty analogous to the analysis of group LASSO (Wainwright, 2019, Chapter 9).
Unlike typical high-dimensional analysis, the result in Theorem 5.1 features two additive terms. The second term arises from the penalty’s null space whose cardinality $|\mathcal{N}|$, understood as the number of roots in the tree (or the forest) structure, is typically much smaller than $n$ and $p$. Hence, the second term is square order for $\sqrt{\log(p)/n}$ and usually becomes negligible with large $n$. Consequently, the first term is more influential in determining the primary error rate: its main driver is
@@ 5.2.1 Direct Comparison with RARE
Consider a setting where the number of true groups in $\beta^{*}$ grows linearly with dimension $p$, so neither $K$ nor $p$ is considered constant in the analysis. In this situation, the tree is made up of $t$ copies of the same subtree, each containing $p_{s}$ parameters within the $K_{s}$ groups. Taking $p_{s}$ and $K_{s}$ as constants, the total dimension $p=t\,p_{s}$ and the total number of groups $K=tK_{s}$ increase with the number of subtree copies $t$. The following corollary shows that, under this scenario, the predictive error rate is at least as good as LASSO’s “slow rate,” matching the rate of Yan and Bien (2021).
@@
Under the expanding tree model in Example 5.1 and the assumptions of Theorem 5.1, with probability at least 12/p1-2/p, we have $n^{-1}\,\|X(\hat{\beta}-\beta^{*})\|_{2}^{2}\;\lesssim\;\sqrt{n^{-1}\log(p)}\,\|\beta^{*}\|_{1},$ where $\lesssim$‘ indicates that we ignore constant factors and higher-order terms.
@@
Suppose that there exists a constant $\rho_{C}$ such that $\rho=\max_{\ell\in\mathcal{B}^{*}}a_{\ell}/\min_{\ell\in\mathcal{B}^{*}}a_{\ell}\leq\rho_{C}$, where $\mathcal{B}^{*}$ is the aggregating set in Definition 3.1 and $\rho_{C}$ is a constant. Then, under the fixed tree model in Example 5.2 the assumptions of Theorem 5.1, with probability at least 12/p1-2/p, $n^{-1}\|X(\hat{\beta}-\beta^{*})\|_{2}^{2}\;\lesssim\;\left[n^{-1/2}+\sqrt{(np)^{-1}\log(p)}\,\right]\|\beta^{*}\|_{1}.$
The result of the preceding corollary suggests an improvement by a factor of $\{\log(p)\}^{-1/2}$ over the method of Yan and Bien (2021), which is further verified by simulation results in Section 6.1.1.